OpenStudy (anonymous):

h(x) = 3x + 3 g(x) = −4x + 1 Find (h + g)(10) Show all work, please :)

hero (hero):

h(x) + g(x) = (3x + 3) + (-4x + 1) = (3x - 4x) + (3 + 1) = -x + 4 = 4 - x (h + g)(10) = 4 - 10

OpenStudy (anonymous):

Thank you.